OpenStudy (anonymous):

Question 1.1. Evaluate the function f(x) = 3x – 2, when x = –1. (Points : 1) –9 –5 1 5

OpenStudy (anonymous):

3(-1)-2=-5

OpenStudy (anonymous):

Yes, when you have f(x) = 3x-2 and it asks you to find x = -1 all you have to do is find f(-1) = 3 (-1) - 2 = -5 as Wendy stated.